# ContextMRI: Enhancing Compressed Sensing MRI through Metadata Conditioning

### v1

@@ Abstract
Compressed sensing MRI seeks to accelerate MRI acquisition processes by sampling fewer $k$-space measurements and then reconstructing the missing data algorithmically. The success of these approaches often relies on strong priors or learned statistical models. While recent diffusion model-based priors have significantly shown great potential, previous methods typically ignore clinically available metadata (e.g. patient demographics, imaging parameters, slice-specific information). In practice, metadata contains meaningful cues about the anatomy and acquisition protocol, suggesting it could further constrain the reconstruction problem. In this work, we propose ContextMRI, a text-conditioned diffusion model for MRI that integrates granular metadata into the reconstruction process. We train a pixel-space diffusion model directly on minimally processed, complex-valued MRI images. During inference, metadata is converted into a structured text prompt and fed to the model via CLIP text embeddings. By conditioning the prior on metadata, we unlock more accurate reconstructions and show consistent gains across multiple datasets, acceleration factors, and undersampling patterns. Our experiments demonstrate that increasing the fidelity of metadata—ranging from slice location and contrast to patient age, sex, and pathology—systematically boosts reconstruction performance. This work highlights the untapped potential of leveraging clinical context for inverse problems and opens a new direction for metadata-driven MRI reconstruction.
{affiliations}
Department of Bio and Brain Engineering, Korea Advanced Institute of Science and Technology, Daejeon, South Korea
Kim Jae Chul Graduate School of AI, Korea Advanced Institute of Science and Technology, Daejeon, South Korea
Department of Computing and Mathematical Sciences, California Institute of Technology, CA, USA
Department of Biomedical Systems Informatics, Yonsei University College of Medicine, Seoul, South Korea
Department of Psychiatry, Yonsei University College of Medicine, Seoul, South Korea
Institute of Behavioral Sciences in Medicine, Yonsei University College of Medicine, Seoul, South Korea
Equal contribution
Correspondence should be addressed to J.C.Y. (jong.ye@kaist.ac.kr)
Abstract
Introduction
Diffusion models [1, 2] have recently emerged as pivotal tools for solving inverse problems in medical imaging [3, 4, 5]. These methods, referred to as diffusion model-based inverse problem solvers (DIS) [6, 7, 8], aim to estimate the posterior distribution of the signal $\bm{x}$ given measurements $\bm{y}$, rather than relying on direct regression from undersampled data to fully sampled reconstructions. By leveraging pre-trained diffusion models as universal priors, DIS approaches can flexibly adapt to different imaging physics (e.g. changes in coil sensitivities, sampling patterns), while providing high-quality reconstruction images in line with Bayesian posterior inference.
@@ Introduction
Figure 2: Quantitative and qualitative comparison of the proposed method against unconditional reconstruction (DDS). The dashed line with $\dagger$ indicates that DDS was performed with the diffusion model used in the original work of [9], which is the reason for the difference in performance between the unconditional version of our method and the dotted line.
@@
Most modern diffusion models trained on natural images are conditional, with text being the most common conditioning  [10, 11, 12]. Training a conditional model provides at least two advantages. First, it enhances controllability with techniques such as classifier-free guidance (CFG) [13]. Second, it was shown that it is significantly easier to train a conditional diffusion model than an unconditional one on the same dataset [14], as text signals can be a useful signal for disentangling the representation space.
@@
In terms of image restoration through DIS, several recent works [15, 16, 17] have shown the promise of leveraging text conditions. It was shown that one can achieve better reconstruction performance by jointly optimizing for the text embeddings [15], and one can use the text condition as the control knob for mode selection [16, 17]. However, the existing methods are very slow to execute [15], require heuristic tuning of the algorithm that is targetted for specific appcliations [16, 17], and often yields over-emphasized results that render the results closer to image editing. Moreover, the situation is inherently different for image restoration and medical image reconstruction. In most image restoration tasks, meaningful text conditioning is scarce because such tasks often lack associated textual metadata that describe the imaging process or the specific content being restored, necessitating algorithms like prompt-tuning to automatically generate useful prompts [15].
@@
In contrast to low-level vision problems explored in [16, 15, 17], in medical imaging, ample metadata exists in various forms. Patient demographics, medical history, body part, imaging parameters of the scan, etc., all provide valuable additional information, which can reshape the prior by conditioning $p(\bm{x}|{\bm{c}})$. These factors have been surprisingly overlooked in the context of inverse problems, where only the measurement signal is used to retrieve the clean signal, a suboptimal practice according to the data processing inequality [18].
@@
Most prior works leveraging text-conditioned generative models for medical data rely on relatively simple and coarse-grained conditioning that is interpretable and often acts closer to discrete class conditioning. Roentgen [19] finetunes Stable Diffusion for X-ray synthesis from simple text descriptions of the pathology. Kim et al. [20] trains a text-to-image LDM for BRATS [21] brain MRI conditioned on contrast and some pathology information. Medisyn [22] trains Würstchen [23], a multi-scale text-to-video LDM on a highly diverse set of data including X-ray, CT, MRI, endoscopy, OCT, etc., but on relatively simpler prompts focused on specifying the modality of the target image. Pinaya et al. [24] trains a 3D LDM that is conditioned on numerical values of age, sex, and ventricular size to generate T1w brain images (UK Biobank [25]). GenerateCT [26] trains a two-stage discrete-continuous text-to-volume diffusion model, which takes in as input age, sex, and impression (i.e. pathology) in text form. The work that is possibly the closest to our work is TUMSyn [27], which is a conditional neural implicit model trained specifically for MRI contrast conversion. The text condition of TUMSyn includes fine-grained information including patient information, scanner type, MR imaging parameters, etc. However, TUMSyn is a model tailored for a specific task of contrast conversion. In contrast, our work focuses on first constructing a general-purpose diffusion generative prior, then using this prior in a zero-shot fashion for solving inverse problems. While we focus on compressed sensing MRI (CS-MRI), our method can also be used for unconditional data synthesis, image inpainting, etc. Another important point to note is that even for text conditional generative models that were trained specifically for the MRI modality [20, 24, 27], none of the works considered the complex-valued nature of MR images, and hence cannot be used for native tasks such as CS-MRI. Moreover, the use of these generative models was mostly constrained to data augmentation [20, 24], or the task that the model was trained for [27]. In what follows, we demonstrate that a pre-trained generative model, which was not specifically trained for any particular MRI inverse problem, can be adapted at inference time to various DIS solvers. This approach achieves a reconstruction quality that surpasses what was previously attainable with unconditional models.
@@
To leverage text conditioning for inverse problems in medical imaging, we propose using a pre-trained generative model that is conditioned on text. Several recent works [19, 28, 24, 20, 22] focus on developing text-to-image diffusion models for image synthesis. However, most of these works focus on improving the generation quality of the images, and the use of the generated data is limited to data augmentation. For the methods that target MRI specifically [24, 20], the models are conditioned by simple text prompts similar to the role of classifiers. A concurrent work of TumSyn [27] trains an implicit neural network for MRI contrast conversion on detailed metadata, but is constrained to a specific task and does not handle the complex-valued nature of MR imaging.
@@
In this work, we propose ContextMRI, the first text-to-image MRI diffusion model that is conditioned on detailed metadata including patient demographics (e.g. age, sex), MR imaging parameters (e.g. TR, TE, etc.), anatomy (e.g. knee, brain), contrast (e.g. PDFS, T2FLAIR), slice location, and pathological findings. Our model is trained on minimum variance unbiased estimate (MVUE) [4] complex-valued images derived from raw $k$-space data and estimated sensitivity maps [29] so that it can be naturally used for downstream tasks such as CS-MRI. Using ContextMRI as a prior, we propose a new paradigm of metadata-conditioned CS-MRI, showing that conditioning the inverse problem-solving process with metadata significantly improves the reconstruction performance, gradually improving with more information. Importantly, we show that ContextMRI consistently outperforms the unconditional counterparts regardless of the circumstances.
Results
Figure 3: Text-guided image synthesis by extracting the metadata from the original image. The text on the bottom shows the extracted text prompt. No measurements from the original images were used in the sampled images.
@@ 0.1 ContextMRI captures intricate metadata cues
A vital requirement for text-conditioned diffusion models in inverse problems is their ability to generate realistic images aligned with the given metadata. Figure 3 demonstrates this alignment qualitatively. We extract metadata from real MRI volumes and feed it into our diffusion model at a CFG scale of 5111CFG scale of 1 is naive conditional inference. Increasing this value emphasizes the conditioning. to emphasize the conditioning signal. Note that no measurement signal from the original image was used in the generation of the images in the second row of Figure 3. Notably, the synthesized images exhibit anatomically coherent structures and contrast weightings consistent with the metadata prompts. For instance, T1POST brain slices display dark cerebrospinal fluid and a bright dural rim, whereas T2FLAIR-like slices depict dark ventricles and sulci against relatively brighter parenchyma. In proton-density (PD) knee scans, bright fat signals, and sharp tissue delineation emerge, while PD fat-suppressed (PDFS) acquisitions suppress fat but preserve fluid hyperintensities. These examples confirm that ContextMRI correctly interprets slice-level metadata, producing anatomically plausible images reflecting the instructed contrast and tissue characteristics.
@@ 0.2 Conditional inference outperforms unconditional inference
We first focus on the case where we have access to all the metadata that was used during training, including anatomy, slice, contrast, and MR imaging parameters. Table 1 reports the quantitative performance of ContextMRI for CS-MRI reconstruction under various acceleration factors, undersampling masks, and anatomies, while Figure2 provides a qualitative summary. We test different CFG scales, from 0.0 (unconditional) to 3.0, and generally see better reconstruction accuracy at higher scales, although in some runs the best performance occurs at lower guidance strengths. We note that when we increase the CFG scale beyond 3.0, the performance starts to drop, which is expected from the known limitations of using high guidance scales [30, 31], especially when coupled with inverse problem solving [32]. This indicates that metadata injection augments the model’s prior beyond information already contained in the undersampled measurement. Furthermore, pushing CFG to around 2–3 provides the best trade-off between leveraging metadata and avoiding over-reliance on text prompts. When pitted against another diffusion model of similar parameter count [^2] Both are U-Net architecture with about 400M parameters. used in the original Decomposed Diffusion Sampler (DDS) framework [9], ContextMRI achieves comparable performance in the unconditional setting. However, once metadata conditioning is enabled, ContextMRI significantly surpasses the DDS baseline. This trend remains consistent regardless of anatomy, mask type, or sampling ratio. Moreover, conditional inference often reduces variability across slices. For example, the structural similarity index measure (SSIM) [33] and the learned perceptual image patch similarity (LPIPS) [34] scores under conditional settings tend to have similar or lower standard deviations compared to the unconditional baseline. This implies that metadata conditioning not only improves average performance but also leads to more stable and consistent reconstructions from slice to slice.
@@ 0.4 Patient Demographics as a Reconstruction Prior
Age-dependent tissue heterogeneity, variations in organ morphology, and progressive degenerative changes all directly influence underlying MR signal characteristics. Likewise, sex-based physiological distinctions—like hormonal fluctuations and structural differences in muscle, bone, and fat distribution—are known to modulate intensity and contrast in MR images. By injecting this demographic context, the diffusion model learns more nuanced priors about the expected image anatomy and contrast patterns. To validate this hypothesis, we fine-tuned our model on SKM-TEA [35] dataset which also contains patient demographic as the metadata, and tested the performance of CS-MRI on the SKM-TEA validation set with uniform1D ×$\times 8$ downsampling. In Fig. 5, we see the reconstruction performance in Peak signal-to-noise-ratio (PSNR) with and without the patient demographic, validating our hypothesis.
@@ 0.5 Pathology Priors and Robustness to Misinformation
Discussion
There has been a recent surge of interest in developing text-to-image diffusion models for medical images, which can potentially be used for diverse clinical practice applications. However, much of the discourse has revolved around synthetic data generation as a primary use case. Synthetic data is often positioned as a generative augmentation tool to improve the performance of downstream tasks, such as segmentation and classification, by enhancing data diversity [19, 26].
Beyond augmentation, the full potential of pre-trained text-to-image diffusion models in medical imaging remains underexplored. In computer vision, these models have demonstrated remarkable versatility, being successfully adapted for tasks like data mining [36], 3D scene reconstruction [37], and solving inverse problems [8], in a zero-shot fashion. The extension of these methodologies to medical imaging could similarly revolutionize clinical workflows. For instance, pre-trained text-to-image models could be re-purposed to generate anatomically consistent counterfactuals for diagnostic reasoning, guide 3D reconstruction of volumetric scans, or act as priors in solving inverse problems such as accelerated MRI reconstruction. Each of these applications could address practical limitations in current clinical practice, paving the way for more integrated, intelligent systems in healthcare.

### v2

@@ Abstract
Compressed sensing MRI seeks to accelerate MRI acquisition processes by sampling fewer $k$-space measurements and then reconstructing the missing data algorithmically. The success of these approaches often relies on strong priors or learned statistical models. While recent diffusion model-based priors have shown great potential, previous methods typically ignore clinically available metadata (e.g. patient demographics, imaging parameters, slice-specific information). In practice, metadata contains meaningful cues about the anatomy and acquisition protocol, suggesting it could further constrain the reconstruction problem. In this work, we propose ContextMRI, a text-conditioned diffusion model for MRI that integrates granular metadata into the reconstruction process. We train a pixel-space diffusion model directly on minimally processed, complex-valued MRI images. During inference, metadata is converted into a structured text prompt and fed to the model via CLIP text embeddings. By conditioning the prior on metadata, we unlock more accurate reconstructions and show consistent gains across multiple datasets, acceleration factors, and undersampling patterns. Our experiments demonstrate that increasing the fidelity of metadata—ranging from slice location and contrast to patient age, sex, and pathology—systematically boosts reconstruction performance. This work highlights the untapped potential of leveraging clinical context for inverse problems and opens a new direction for metadata-driven MRI reconstruction. Code is available at https://github.com/DoHunLee1/ContextMRI
@@
Figure 2: Quantitative and qualitative comparison of the proposed method against unconditional reconstruction (DDS). The dashed line with $\dagger$ indicates that DDS was performed with the diffusion model used in the original work of [chung2024decomposed], which is the reason for the difference in performance between the unconditional version of our method and the dotted line.
1 Introduction
Diffusion models [song2020score, ho2020denoising] have recently emerged as pivotal tools for solving inverse problems in medical imaging [chung2022score, jalal2021robust, song2022solving]. These methods, referred to as diffusion model-based inverse problem solvers (DIS) [kadkhodaie2021stochastic, kawar2022denoising, chung2023diffusion], aim to estimate the posterior distribution of the signal ${\bm{x}}$ given measurements ${\bm{y}}$, rather than relying on direct regression from undersampled data to fully sampled reconstructions. By leveraging pre-trained diffusion models as universal priors, DIS approaches can flexibly adapt to different imaging physics (e.g. changes in coil sensitivities, sampling patterns), while providing high-quality reconstruction images in line with Bayesian posterior inference.
@@ 1 Introduction
Most modern diffusion models trained on natural images are conditional, with text being the most common conditioning  [rombach2022high, saharia2022photorealistic, videoworldsimulators2024]. Training a conditional model provides at least two advantages. First, it enhances controllability with techniques such as classifier-free guidance (CFG) [ho2021classifierfree]. Second, it was shown that it is significantly easier to train a conditional diffusion model than an unconditional one on the same dataset [peebles2023scalable], as text signals can be a useful signal for disentangling the representation space.
@@
In terms of image restoration through DIS, several recent works [chung2024prompttuning, kim2023regularization, kim2025dreamsampler] have shown the promise of leveraging text conditions. It was shown that one can achieve better reconstruction performance by jointly optimizing for the text embeddings [chung2024prompttuning], and one can use the text condition as the control knob for mode selection [kim2023regularization, kim2025dreamsampler]. However, the existing methods are very slow to execute [chung2024prompttuning], require heuristic tuning of the algorithm that is targetted for specific appcliations [kim2023regularization, kim2025dreamsampler], and often yields over-emphasized results that render the results closer to image editing. Moreover, the situation is inherently different for image restoration and medical image reconstruction. In most image restoration tasks, meaningful text conditioning is scarce because such tasks often lack associated textual metadata that describe the imaging process or the specific content being restored, necessitating algorithms like prompt-tuning to automatically generate useful prompts [chung2024prompttuning].
@@
In contrast to low-level vision problems explored in [kim2023regularization, chung2024prompttuning, kim2025dreamsampler], in medical imaging, ample metadata exists in various forms. Patient demographics, medical history, body part, imaging parameters of the scan, etc., all provide valuable additional information, which can reshape the prior by conditioning $p({\bm{x}}|{\bm{c}})$. These factors have been surprisingly overlooked in the context of inverse problems, where only the measurement signal is used to retrieve the clean signal, a suboptimal practice according to the data processing inequality [cover1999elements].
@@
Most prior works leveraging text-conditioned generative models for medical data rely on relatively simple and coarse-grained conditioning that is interpretable and often acts closer to discrete class conditioning. Roentgen [bluethgen2024vision] finetunes Stable Diffusion for X-ray synthesis from simple text descriptions of the pathology. Kim et al. [kim2024controllable] trains a text-to-image LDM for BRATS [menze2014multimodal] brain MRI conditioned on contrast and some pathology information. Medisyn [cho2024medisyn] trains Würstchen [pernias2023wurstchen], a multi-scale text-to-video LDM on a highly diverse set of data including X-ray, CT, MRI, endoscopy, OCT, etc., but on relatively simpler prompts focused on specifying the modality of the target image. Pinaya et al. [pinaya2022brain] trains a 3D LDM that is conditioned on numerical values of age, sex, and ventricular size to generate T1w brain images (UK Biobank [sudlow2015uk]). GenerateCT [hamamci2023generatect] trains a two-stage discrete-continuous text-to-volume diffusion model, which takes in as input age, sex, and impression (i.e. pathology) in text form. The work that is possibly the closest to our work is TUMSyn [wang2024towards], which is a conditional neural implicit model trained specifically for MRI contrast conversion. The text condition of TUMSyn includes fine-grained information including patient information, scanner type, MR imaging parameters, etc. However, TUMSyn is a model tailored for a specific task of contrast conversion. In contrast, our work focuses on first constructing a general-purpose diffusion generative prior, then using this prior in a zero-shot fashion for solving inverse problems. While we focus on compressed sensing MRI (CS-MRI), our method can also be used for unconditional data synthesis, image inpainting, etc. Another important point to note is that even for text conditional generative models that were trained specifically for the MRI modality [kim2024controllable, pinaya2022brain, wang2024towards], none of the works considered the complex-valued nature of MR images, and hence cannot be used for native tasks such as CS-MRI. Moreover, the use of these generative models was mostly constrained to data augmentation [kim2024controllable, pinaya2022brain], or the task that the model was trained for [wang2024towards]. In what follows, we demonstrate that a pre-trained generative model, which was not specifically trained for any particular MRI inverse problem, can be adapted at inference time to various DIS solvers. This approach achieves a reconstruction quality that surpasses what was previously attainable with unconditional models.
@@
To leverage text conditioning for inverse problems in medical imaging, we propose using a pre-trained generative model that is conditioned on text. Several recent works [bluethgen2024vision, khader2023denoising, pinaya2022brain, kim2024controllable, cho2024medisyn] focus on developing text-to-image diffusion models for image synthesis. However, most of these works focus on improving the generation quality of the images, and the use of the generated data is limited to data augmentation. For the methods that target MRI specifically [pinaya2022brain, kim2024controllable], the models are conditioned by simple text prompts similar to the role of classifiers. A concurrent work of TumSyn [wang2024towards] trains an implicit neural network for MRI contrast conversion on detailed metadata, but is constrained to a specific task and does not handle the complex-valued nature of MR imaging.
@@
In this work, we propose ContextMRI, the first text-to-image MRI diffusion model that is conditioned on detailed metadata including patient demographics (e.g. age, sex), MR imaging parameters (e.g. TR, TE, etc.), anatomy (e.g. knee, brain), contrast (e.g. PDFS, T2FLAIR), slice location, and pathological findings. Our model is trained on minimum variance unbiased estimate (MVUE) [jalal2021robust] complex-valued images derived from raw $k$-space data and estimated sensitivity maps [uecker2014espirit] so that it can be naturally used for downstream tasks such as CS-MRI. Using ContextMRI as a prior, we propose a new paradigm of metadata-conditioned CS-MRI, showing that conditioning the inverse problem-solving process with metadata significantly improves the reconstruction performance, gradually improving with more information. Importantly, we show that ContextMRI consistently outperforms the unconditional counterparts regardless of the circumstances.
2 Result
@@ 2.1 ContextMRI captures intricate metadata cues
A vital requirement for text-conditioned diffusion models in inverse problems is their ability to generate realistic images aligned with the given metadata. Figure LABEL:fig:image_synthesis demonstrates this alignment qualitatively. We extract metadata from real MRI volumes and feed it into our diffusion model at a CFG scale of 5111CFG scale of 1 is naive conditional inference. Increasing this value emphasizes the conditioning. to emphasize the conditioning signal. Note that no measurement signal from the original image was used in the generation of the images in the second row of Figure LABEL:fig:image_synthesis. Notably, the synthesized images exhibit anatomically coherent structures and contrast weightings consistent with the metadata prompts. For instance, T1POST brain slices display dark cerebrospinal fluid and a bright dural rim, whereas T2FLAIR-like slices depict dark ventricles and sulci against relatively brighter parenchyma. In proton-density (PD) knee scans, bright fat signals, and sharp tissue delineation emerge, while PD fat-suppressed (PDFS) acquisitions suppress fat but preserve fluid hyperintensities. These examples confirm that ContextMRI correctly interprets slice-level metadata, producing anatomically plausible images reflecting the instructed contrast and tissue characteristics.
@@ 2.2 Conditional inference outperforms unconditional inference
We first focus on the case where we have access to all the metadata that was used during training, including anatomy, slice, contrast, and MR imaging parameters. Table 1 reports the quantitative performance of ContextMRI for CS-MRI reconstruction under various acceleration factors, undersampling masks, and anatomies, while Figure2 provides a qualitative summary. We test different CFG scales, from 0.0 (unconditional) to 3.0, and generally see better reconstruction accuracy at higher scales, although in some runs the best performance occurs at lower guidance strengths. We note that when we increase the CFG scale beyond 3.0, the performance starts to drop, which is expected from the known limitations of using high guidance scales [poole2022dreamfusion, bradley2024classifier], especially when coupled with inverse problem solving [chung2024cfg++]. This indicates that metadata injection augments the model’s prior beyond information already contained in the undersampled measurement. Furthermore, pushing CFG to around 2–3 provides the best trade-off between leveraging metadata and avoiding over-reliance on text prompts. When pitted against another diffusion model of similar parameter count [^2] Both are U-Net architecture with about 400M parameters. used in the original Decomposed Diffusion Sampler (DDS) framework [chung2024decomposed], ContextMRI achieves comparable performance in the unconditional setting. However, once metadata conditioning is enabled, ContextMRI significantly surpasses the DDS baseline. This trend remains consistent regardless of anatomy, mask type, or sampling ratio. Moreover, conditional inference often reduces variability across slices. For example, the structural similarity index measure (SSIM) [Wang2004] and the learned perceptual image patch similarity (LPIPS) [zhang2018unreasonable] scores under conditional settings tend to have similar or lower standard deviations compared to the unconditional baseline. This implies that metadata conditioning not only improves average performance but also leads to more stable and consistent reconstructions from slice to slice.
@@ 2.4 Patient Demographics as a Reconstruction Prior
Age-dependent tissue heterogeneity, variations in organ morphology, and progressive degenerative changes all directly influence underlying MR signal characteristics. Likewise, sex-based physiological distinctions—like hormonal fluctuations and structural differences in muscle, bone, and fat distribution—are known to modulate intensity and contrast in MR images. By injecting this demographic context, the diffusion model learns more nuanced priors about the expected image anatomy and contrast patterns. To validate this hypothesis, we fine-tuned our model on SKM-TEA [desai2022skm] dataset which also contains patient demographic as the metadata, and tested the performance of CS-MRI on the SKM-TEA validation set with uniform1D ×$\times 8$ downsampling. In Fig. 4, we see the reconstruction performance in Peak signal-to-noise-ratio (PSNR) with and without the patient demographic, validating our hypothesis.
@@ 2.5 Pathology Priors and Robustness to Misinformation
3 Discussion
There has been a recent surge of interest in developing text-to-image diffusion models for medical images, which can potentially be used for diverse clinical practice applications. However, much of the discourse has revolved around synthetic data generation as a primary use case. Synthetic data is often positioned as a generative augmentation tool to improve the performance of downstream tasks, such as segmentation and classification, by enhancing data diversity [bluethgen2024vision, hamamci2023generatect].
Beyond augmentation, the full potential of pre-trained text-to-image diffusion models in medical imaging remains underexplored. In computer vision, these models have demonstrated remarkable versatility, being successfully adapted for tasks like data mining [siglidis2024diffusion], 3D scene reconstruction [chung2023luciddreamer], and solving inverse problems [chung2023diffusion], in a zero-shot fashion. The extension of these methodologies to medical imaging could similarly revolutionize clinical workflows. For instance, pre-trained text-to-image models could be re-purposed to generate anatomically consistent counterfactuals for diagnostic reasoning, guide 3D reconstruction of volumetric scans, or act as priors in solving inverse problems such as accelerated MRI reconstruction. Each of these applications could address practical limitations in current clinical practice, paving the way for more integrated, intelligent systems in healthcare.